OpenStudy (anonymous):

y'- y = t^2 solution of the equation as a power series??

OpenStudy (amistre64):

define y and y' in terms of power series, what do we get?

OpenStudy (amistre64):

\[y=\sum_0 c_n~t^n\] \[y'=\sum_1 c_n~nt^{n-1}\] use them in the setup \[\sum_0 c_n~nt^{n-1}~-~\sum_1 c_n~t^n=t^2\] need to align exponents and indexes

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

had my 0 and 1 under the sums a bit off track lol \[\sum_{1} c_n~nt^{n-1}~-~\sum_{0+1} c_{n-1}(n-1)~t^{n-1}=t^2\] \[\sum_{1} c_n~nt^{n-1}~-~\sum_{1} c_{n-1}(n-1)~t^{n-1}=t^2\] now we have it so that the indexes and exponents line up, we can combine the summations

OpenStudy (amistre64):

\[\sum_{1} [c_n- c_{n-1}(n-1)]~t^{n-1}=t^2\] lets start poping out terms \[c_1+\sum_{2} [c_n- c_{n-1}(n-1)]~t^{n-1}=t^2\] \[c_1+[c_2- c_{1}(1)]~t+\sum_{3} [c_n- c_{n-1}(n-1)]~t^{n-1}=t^2\] \[c_1+[c_2- c_{1}(1)]~t+[c_3- c_{2}(2)]~t^2+\sum_{4} [c_n- c_{n-1}(n-1)]~t^{n-1}=t^2\]

OpenStudy (amistre64):

comparing coefficients: \[c_1=0\\ c_2-c_1=0\\ c_3-2c_2=1\] and all the others equal 0

OpenStudy (amistre64):

c2=c1=0, c3=1

OpenStudy (amistre64):

and for n>3\[c_n- c_{n-1}(n-1)=0\] \[c_n= (n-1)c_{n-1}\] c1 = 0 c2 = 0 c3 = 1 c4 = 3 c3 = 3 c5 = 4 c4 = 12 c6 = 5 c5 = 60 etc ...

OpenStudy (amistre64):

not real sure how we could develop a rule for the sequence of coeffs tho :/

OpenStudy (amistre64):

i spose we could say: y = t^2 + summation formula

OpenStudy (amistre64):

any ideas if this has an error in it?

OpenStudy (anonymous):

development is c1 + c2.t^2 + c3 t^3...

OpenStudy (anonymous):

I think it is correct what you did

OpenStudy (amistre64):

c0 + c1 t + c2 t^2 + c3 t^3 etc .... such that everything but c2 is zero, and c2 = 1 im outta practice to determine if im missing something along the way :/

OpenStudy (amistre64):

do we account for the homogenous solution first? then assess it for the particulars?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

cn - (n-1)c{n-1} = 0 cn = (n-1)c{n-1} c1 = 0 c0 c2 = 1*0 c0 c3 = 2*1*0 c0 im gonna need a review :) and i gotta go for tonight